mathbuddy:

QUESTION 1 (08.06A) The equation of the line of best fit of a scatter plot is y = 12x + 7. What is the slope of the equation? (4 points) -12 -7 12 7 4 points QUESTION 2 (08.06A) The equation of the line of best fit of a scatter plot is y = −7x − 2. What is the the y-intercept? (4 points) -7 -2 2 7

mathbuddy:

can some help me understand these?

ScaryPerson:

maybe post a ss.

mathbuddy:

?

ScaryPerson:

post a screen shot of the question

mathbuddy:

oh i on a chrome os i don't know how to screenshot

ScaryPerson:

ok i got you.

ScaryPerson:

Press Shift + Ctrl + that button at the top of the screen that had a square with 2 lines next to it

carmelle:

Hi there! I can help you!! The equation for slope-intercept is `y=mx+b` where `m = slope` and `y = y-intercept` Lets start with question 1, `y = 12x + 7` what do u think the slope is?

mathbuddy:

ok back

mathbuddy:

oh and 12

carmelle:

Thats correct! Now for question 2, `y = −7x − 2` whats the y-intercept? remember to pay attention to the signs!!

mathbuddy:

-7

mathbuddy:

?

carmelle:

no that wouldn't be correct y=mx+b where m = slope and b = y-intercept sorry, I wrote this but there was a typo the equation: y = −7x − 2 m = -7 b = ?

mathbuddy:

oh -2

carmelle:

Yep, that would be correct. Sorry about the typo, I didn't mean to confuse u :/

mathbuddy:

it ok

mathbuddy:

thx for the help

carmelle:

Anytime!